OpenStudy (anonymous):

Solve.....(x-3)(x+3) = 6x-14

OpenStudy (king):

x=5,1

OpenStudy (hoblos):

=> x^2 - 9 -6x +14 =0 => x^2 -6x + 5 =0 => (x-1)(x-5)=0 => x=1 or x=5

OpenStudy (king):

here also simplify and use quadratic equation formula